OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1. 4•6+5•7+6•8+...+4n(4n+2)=(4(4n+1)(8n+7)) / (6)

OpenStudy (anonymous):

so you know that mathematical induction has four important steps to follow, right?

OpenStudy (anonymous):

no please explain

OpenStudy (anonymous):

@Percy* @BAdhi @thomas5267

OpenStudy (thomas5267):

How is this even right? Take n=1 LHS=4x6=24 RHS=4x5x15/6=300/6=50

OpenStudy (thomas5267):

I think the statement is incorrect.

OpenStudy (anonymous):

yeah is incorrect but you need to mention that, but the principles of mathematical induction, 1. 4•6+5•7+6•8+...+4n(4n+2)=(4(4n+1)(8n+7)) / (6) is not true for all integer n

OpenStudy (anonymous):

it just has ? marks everywhere on your response @Percy*

OpenStudy (anonymous):

and can y'all help me with one more?

OpenStudy (thomas5267):

You don't even need the principle of mathematical induction. If LHS does not equal RHS for one n, then the whole thing is invalid.

OpenStudy (anonymous):

sorry, i dont know where this ? is coming from but 4get ?, coz look the equation... thomas i think we do becasue the question ask us to

OpenStudy (anonymous):

Here is my second one, you do the same thing. 1^2 + 4^2 + 7^2 + ... + (3n-2)^2 = (n(6n^2 - 3n - 1)) / 2

OpenStudy (anonymous):

yeah, but what is the restriction say

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

remember that our restriction for the first one says for all positive integers, that gave us a clue where to start

OpenStudy (anonymous):

the second one has the exact same instructions.

OpenStudy (anonymous):

yeah, you do the same thing

OpenStudy (anonymous):

will you type it out please

OpenStudy (anonymous):

or show me what i'm supposed to write

OpenStudy (anonymous):

oky, this will b long, coz we will do step by step

OpenStudy (anonymous):

ok but last time your response had a lot of ?'s in it because I guess you used symbols that my computer wasnt familiar with or something

OpenStudy (anonymous):

no dont worry

OpenStudy (anonymous):

so we have a basis step, inductive hypothesis, inductive step and conclusions

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

guys i just typed somthing and doesnt show up....

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

let me try again

OpenStudy (anonymous):

basis step( let n=1)\[1^2+4^2+7^2+...(3n-2)=\frac{ n(6n^2-3n-1) }{ 2}\] \[lhs=(3n-2)^2=(3(1)-2)^2\]=1\[rhs=\frac{ n(6n^2-3n-1) }{ 2 }=\frac{ 1(6(1)-3(1)-1) }{ 2}=1\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

therefore LHS=RHS, do you see dat

OpenStudy (anonymous):

yes

OpenStudy (thomas5267):

\[\text{Use mathematical induction to prove the statement is true}\\ \text{for all positive integers n, or }\textit{show why it is false.}\]

OpenStudy (thomas5267):

The statement is not necessarily true.

OpenStudy (anonymous):

the thing is we not done

OpenStudy (anonymous):

the inductive hypothesis: for k, a fixed but arbitrary integer with k>=1, assume that results holds when n=k, such that:\[1^2+4^2...(3k-2)^2=\frac{ k(6k^2-3k-1) }{ 2 }\]

OpenStudy (anonymous):

do you see that